# Islands in Kerr-Newman Black Holes

### v1

@@ Abstract
We investigate the information paradox in the four-dimensional Kerr-Newman black hole by employing the recently proposed island paradigm. We first consider the quantum filed in the four-dimensional Kerr-Newman spacetime. By employing the near-horizon limit, we demonstrate that the field can be effectively described by a reduced two-dimensional field. Consequently, the formula of entanglement entropy in CFT2 can be naturally adapted to this reduced two-dimensional theory. Under the framework of this reduced two-dimensional theory, we show that the entanglement entropy of radiation for the non-extremal case satisfies the unitarity in the later stage of the appearance of the entanglement islands. We further examine the impact of angular momentum and charges on the Page time and the scrambling time. Both quantities increases as the angular momentum increases, while decreases as the charge increases. At last, we consider the near extremal case. Resort to the Kerr/CFT correspondence, the near-horizon geometry of near extremal Kerr-Newman black holes can be taken account for a warped AdS geometry. In this scenario, the low-energy effective degrees of freedom are dominated by the Schwarzian zero mode, resulting in a one-loop correction to the partition function. The entanglement entropy is subsequently recalculated under the thermodynamic with corrections. Through explicit calculations, we finally find that the Page time and the scrambling time exhibits quantum delays. This strongly suggests that the near extremal geometry is governed by the Schwarzschild dynamics, in which quantum fluctuations result in a reduced rate of information leakage. Our findings further substantiate the conservation of information and extend the applicability of the island paradigm to the most general stationary spacetime background.
@@ I Introduction
At present, the island paradigm has not only be applied to the initial evaporative Jackiw-Teitelboim gravity jtgravity1 ; jtgravity2 but also extends to scenarios involving eternal black holes eternalbh and various specific spacetime background jtisland ; 2d1 ; 2d2 ; 2d3 ; 2d4 ; 2d5 ; 2d6 ; ymh2d ; ymhjt ; 2d7 ; ds1 ; ds2 ; ds3 ; ds4 ; ds5 ; cosmology1 ; cosmology2 ; cosmology3 ; cosmology4 ; cosmology5 ; cosmology6 ; wedgeholo1 ; wedgeholo2 ; wedgeholo3 ; wedgeholo4 ; wedgeholo5 ; wedgeholo6 ; negative ; reflectedentropy1 ; reflectedentropy2 ; other1 ; other2 ; other3 ; other4 ; other5 ; other6 ; sch ; rn ; kerr ; kerrds ; extremalbh1 ; extremalbh2 ; other7 ; other8 ; other9 ; other10 ; other11 ; other12 ; other13 ; other14 ; other15 ; other16 ; other17 ; other18 ; other19 ; ymhghs ; ymhbtz ; ymhacc ; ymhrn ; ymhqfc ; ymhmodular ; other20 ; other21 ; other22 ; other23 ; other24 ; other25 ; other26 ; other27 ; other28 ; other29 ; other30 ; other31 ; other32 ; other33 ; other34 ; other35 ; other36 ; other37 ; other38 ; other39 ; other40 ; other41 . However, in these studies, nearly all have focused on the case of black holes that are static spherically symmetric jtisland ; 2d1 ; 2d2 ; 2d3 ; 2d4 ; 2d5 ; 2d6 ; ymh2d ; ymhjt ; 2d7 ; other1 ; other2 ; other3 ; other4 ; other5 ; other6 ; sch ; rn ; kerr ; kerrds ; extremalbh1 ; extremalbh2 ; other7 ; other8 ; other9 ; other10 ; other11 ; other12 ; other13 ; other14 ; other15 ; other16 ; other17 ; other18 ; other19 ; ymhghs ; ymhbtz ; ymhacc ; ymhrn ; ymhqfc ; ymhmodular ; other20 ; other21 ; other22 ; other23 ; other24 ; other25 ; other26 ; other27 ; other28 ; other29 ; other30 ; other31 ; other32 ; other33 ; other34 ; other35 ; other36 ; other37 ; other38 ; other39 ; other40 . There are few reports on non-spherically symmetric stationary black holes ymhbtz ; kerr ; kerrds . For the most general type in four-dimensional spacetime, known as the Kerr-Newman black hole. Its metric describes a rotating, charged mass and represents the most general solution to the Einstein’s equations in GR. Therefore, it holds significant theoretical importance in the mathematical framework of GR and extends beyond. Correspondingly, resolving the information paradox in this black hole is a critical issue. This study can also offer valuable insights into other four-dimensional scenarios.
@@
The structure of this paper is organized as follows. In section II, we briefly review the fundamental properties of the Kerr-Newman black hole and introduce the standard technique that can transform a quantum field in the four-dimensional Kerr-Newman spacetime into an effective two-dimensional description. This approach establishes an effective two-dimensional metric near the event horizon, which facilitates the subsequent calculation of the entanglement entropy. In section III, we rigorously calculate the entanglement entropy of the radiation emitted by non-extremal Kerr-Newman black holes and derive the corresponding Page curve. Furthermore, we analyze the effects of the charge and the angular momentum on both the Page time and the scrambling time. In section IV, we focus on the near extremal cases. Based on the Kerr/CFT correspondence, the near extremal Kerr-Newman geometry approximates to a warped AdS3 at the near-horizon region. By incorporating the one-loop correction to the modified thermodynamics through the zero mode associated with Schwarzian dynamics, we arrive at a final result of Page time and scrambling time. We find these quantities will be significantly delayed due to the one-loop correction and provide the correct result in the extremal limit. Finally, the section V summarizes the conclusions and provides further discussions.

### v2

@@ Abstract
We investigate the information paradox in the four-dimensional Kerr-Newman black hole by employing the recently proposed island paradigm. We first consider the quantum field in the four-dimensional Kerr-Newman spacetime. By employing the near-horizon limit, we demonstrate that the field can be effectively described by a reduced two-dimensional field theory. Consequently, the formula of entanglement entropy in CFT2 can be naturally adapted to this reduced two-dimensional theory. Under the framework of this reduced two-dimensional theory, we show that the entanglement entropy of radiation for the non-extremal case satisfies the unitarity in the later stage of the appearance of the entanglement islands. We further examine the impact of angular momentum and charges on the Page time and the scrambling time. Both quantities increases as the angular momentum increases, while decreases as the charge increases. At last, we consider the near extremal case. Resort to the Kerr/CFT correspondence, the near-horizon geometry of near extremal Kerr-Newman black holes can be taken account for a warped AdS geometry. In this scenario, the low-energy effective degrees of freedom are dominated by the Schwarzian zero mode, resulting in a one-loop correction to the partition function. The entanglement entropy is subsequently recalculated under the thermodynamic with corrections. Through explicit calculations, we finally find that the Page time and the scrambling time exhibits quantum delays. This strongly suggests that the near extremal geometry is governed by the Schwarzian dynamics, in which quantum fluctuations result in a reduced rate of information leakage. Our findings further substantiate the conservation of information and extend the applicability of the island paradigm to the most general stationary spacetime background.
@@ I Introduction
At present, the island paradigm has not only be applied to the initial evaporative Jackiw-Teitelboim gravity jtgravity1 ; jtgravity2 but also extends to scenarios involving eternal black holes eternalbh and various specific spacetime background jtisland ; 2d1 ; 2d2 ; 2d3 ; 2d4 ; 2d5 ; 2d6 ; ymh2d ; ymhjt ; 2d7 ; ds1 ; ds2 ; ds3 ; ds4 ; ds5 ; cosmology1 ; cosmology2 ; cosmology3 ; cosmology4 ; cosmology5 ; cosmology6 ; cosmology7 ; wedgeholo1 ; wedgeholo2 ; wedgeholo3 ; wedgeholo4 ; wedgeholo5 ; wedgeholo6 ; negative ; reflectedentropy1 ; reflectedentropy2 ; other1 ; other2 ; other3 ; other4 ; other5 ; other6 ; sch ; rn ; kerr ; kerrds ; extremalbh1 ; extremalbh2 ; other7 ; other8 ; other9 ; other10 ; other11 ; other12 ; other13 ; other14 ; other15 ; other16 ; other17 ; other18 ; other19 ; ymhghs ; ymhbtz ; ymhacc ; ymhrn ; ymhqfc ; ymhmodular ; other20 ; other21 ; other22 ; other23 ; other24 ; other25 ; other26 ; other27 ; other28 ; other29 ; other30 ; other31 ; other32 ; other33 ; other34 ; other35 ; other36 ; other37 ; other38 ; other39 ; other40 ; other41 ; other42 ; other43 ; other44 ; other45 ; other46 ; other47 ; other48 . However, in these studies, nearly all have focused on the case of black holes that are static spherically symmetric jtisland ; 2d1 ; 2d2 ; 2d3 ; 2d4 ; 2d5 ; 2d6 ; ymh2d ; ymhjt ; 2d7 ; other1 ; other2 ; other3 ; other4 ; other5 ; other6 ; sch ; rn ; kerr ; kerrds ; extremalbh1 ; extremalbh2 ; other7 ; other8 ; other9 ; other10 ; other11 ; other12 ; other13 ; other14 ; other15 ; other16 ; other17 ; other18 ; other19 ; ymhghs ; ymhbtz ; ymhacc ; ymhrn ; ymhqfc ; ymhmodular ; other20 ; other21 ; other22 ; other23 ; other24 ; other25 ; other26 ; other27 ; other28 ; other29 ; other30 ; other31 ; other32 ; other33 ; other34 ; other35 ; other36 ; other37 ; other38 ; other39 ; other40 . There are few reports on non-spherically symmetric stationary black holes ymhbtz ; kerr ; kerrds . For the most general type in four-dimensional spacetime, known as the Kerr-Newman black hole. Its metric describes a rotating, charged mass and represents the most general solution to the Einstein’s equations in GR. Therefore, it holds significant theoretical importance in the mathematical framework of GR and extends beyond. Correspondingly, resolving the information paradox in this black hole is a critical issue. This study can also offer valuable insights into other four-dimensional scenarios.
@@
The structure of this paper is organized as follows. In section II, we briefly review the fundamental properties of the Kerr-Newman black hole and introduce the standard technique that can transform a quantum field in the four-dimensional Kerr-Newman spacetime into an effective two-dimensional description. This approach establishes an effective two-dimensional metric near the event horizon, which facilitates the subsequent calculation of the entanglement entropy. In section III, we rigorously calculate the entanglement entropy of the radiation emitted by non-extremal Kerr-Newman black holes and derive the corresponding Page curve. Furthermore, we analyze the effects of the charge and the angular momentum on both the Page time and the scrambling time. In section IV, we focus on the near extremal cases. Based on the Kerr/CFT correspondence, the near extremal Kerr-Newman geometry approximates to a warped AdS3 at the near-horizon region. By incorporating the one-loop correction to the modified thermodynamics through the zero mode associated with Schwarzian dynamics, we arrive at a final result of Page time and scrambling time. We find these quantities will be significantly delayed due to the one-loop correction and provide the correct result in the near extremal limit. Finally, the section V summarizes the conclusions and provides further discussions.